If Sin2A=Sin3A, then value of A is:
A. 18
B. 36
C. 54
D. 72

Grade:12th pass

2 Answers

Aditya Gupta
4 years ago
hello pushpa. this ques can be easily solved as follows
sin2A= sin3A
2A= 3A or A= 0 (not given in options).
so we try another approach
we use the formula sinx= sin(180 – x)
so, sin2A= sin3A= sin(180 – 3A)
so 2A= 180 – 3A
or A= 180/5
or A= 36 deg
option B
kindly approve :)
Saurabh Koranglekar
4 years ago
Dear student

Sin ( 90- x ) = sin ( 90 + x)

2a= 90-x

3a = 90+x

5a = 180

a= 36

Regards